The total revenue function for a certain product is given by The total revenue function for a certain product is given by R=630x dollars, and the total cost function for this product is
Answer #1

@ Given revenue function R = 630x and cost function Cz 10,000+ 30x+x2 a profit function P=R-C -)P: 630x - [10,000 +30% +x2] p- 2n- boo x = 600 2300 12=300 Again differentiate eg (2) wir to x adap lad (600-2x) duldu, dn dap d 600-224 du? In de d4 =0-2